Phantomdex:

A principal amount of $5,000 will be placed in a savings account. A simple interest account earns an annual percentage rate of 6.8%, while a compound interest account has an APR of 3.7% compounded quarterly. Which account has a larger balance after 20 years? The compound interest account has a larger balance of $11,800. The simple interest account has a larger balance of $10,444.09. The simple interest account has a larger balance of $11,800. The compound interest account has a larger balance of $10,444.09.

SmoothCriminal:

Ay dios mío

SmoothCriminal:

Give me a sec, this is a gigantic interest amount for SO much time.

SmoothCriminal:

\begin{enumerate} \item The compound interest account has a larger balance of $11,800. \item The simple interest account has a larger balance of $10,444.09. \item The simple interest account has a larger balance of $11,800. \item The compound interest account has a larger balance of $10,444.09. \end{enumerate} Now to compare the two savings accounts, you can use the formuIas for simple interest and compound interest: For simple interest, the formuIa I'm using is: \[ A_{\text{simple}} = P \left(1 + rt\right) \] This is for "Where": \(A_{\text{simple}}\) = the future value of the investment, including interest \(P\) = the principal investment amount = $5,000 \(r\) = the annual interest rate (decimal) = 6.8% or 0.068 \(t\) = the time the money is invested for, in years = 20 years Now, you'll calculate the future value for the simple interest account like this 👇 \[ A_{\text{simple}} = 5,000 \left(1 + 0.068 \cdot 20\right) = 5,000 \cdot 2.36 = \$11,800 \] Now that we got that crap out the way, we'll move to the compound interest 🤏🤓 For compound interest, the formula I'm using is: \[ A_{\text{compound}} = P \left(1 + \frac{r}{n}\right)^{nt} \] Where (same thing as before): \(A_{\text{compound}}\) = the future value of the investment, including interest \(P\) = the principal investment amount = $5,000 \(r\) = the annual interest rate (decimal) = 3.7% or 0.037 \(n\) = the number of times that interest is compounded per year = 4 (quarterly) \(t\) = the time the money is invested for, in years = 20 years Now finally you'll calculate the future value for the compound interest account: \[ A_{\text{compound}} = 5,000 \left(1 + \frac{0.037}{4}\right)^{4 \cdot 20} \approx 5,000 \cdot 1.2391 \approx \$6,195.47 \] Then cmparing the two future values 👇 The compound interest account has a larger balance of $6,195.47 compared to the regular interest account's balance of $11,800. Now the summarized answer is down below 👇 The simple interest account has a larger balance of $11,800. 👌 😎

SmoothCriminal:

Comparing*

Phantomdex:

Great, I have more questions-

SmoothCriminal:

@phantomdex wrote:
Great, I have more questions-
Amazing! Let's do it.

Phantomdex:

Just gonna post it here because I'm lazy

Phantomdex:

A principal amount of $400 is placed in a savings account with an annual rate of 5.75% simple interest for 8 years. Which formula would correctly determine the final amount? A = (400)(0.0575)(8) A = (400)(5.75)(8) A = (400)(5.75)(8) + 400 A = (400)(0.0575)(8) + 400

SmoothCriminal:

@phantomdex wrote:
A principal amount of $400 is placed in a savings account with an annual rate of 5.75% simple interest for 8 years. Which formula would correctly determine the final amount? A = (400)(0.0575)(8) A = (400)(5.75)(8) A = (400)(5.75)(8) + 400 A = (400)(0.0575)(8) + 400
Finally an easy one

SmoothCriminal:

@phantomdex wrote:
A principal amount of $400 is placed in a savings account with an annual rate of 5.75% simple interest for 8 years. Which formula would correctly determine the final amount? A = (400)(0.0575)(8) A = (400)(5.75)(8) A = (400)(5.75)(8) + 400 A = (400)(0.0575)(8) + 400
So the W formula to determine the amount with simple interest is: 👇 \[A = (400)(0.0575)(8) + 400\]

Phantomdex:

A principal amount of $900 will be placed in a savings account for 13 years with an APR of 2.9%. What is the difference in the total account balances if simple interest is applied, compared to monthly compounded interest? $972.22 $339.30 $72.22 $27.22

SmoothCriminal:

@phantomdex wrote:
A principal amount of $900 will be placed in a savings account for 13 years with an APR of 2.9%. What is the difference in the total account balances if simple interest is applied, compared to monthly compounded interest? $972.22 $339.30 $72.22 $27.22
Damn, alright. Good thing I'm doing math homework right now. First and foremost, use my formulas. They never fail 😏 Here (they always work) 💪: For simple interest: \[A_{\text{simple}} = P \cdot (1 + rt)\] For monthly compounded interest: \[A_{\text{monthly}} = P \cdot \left(1 + \frac{r}{n}\right)^{nt}\] Where: - \(A_{\text{simple}}\) is the total account balance with simple interest. - \(A_{\text{monthly}}\) is the total account balance with monthly compounded interest. - \(P\) is the principal magnitude, which is $900 in this case. - \(r\) is the annual interest rate (decimal), which is 2.9% or 0.029. - \(t\) is the time the mazuma is invested for, which is 13 years. - \(n\) is the number of times interest is compounded per year, which is 12 for monthly compounding. Now, let's calculate both account balances: For simple interest: \[A_{\text{simple}} = 900 \cdot \left(1 + 0.029 \cdot 13\right) = 900 \cdot 1.377 = $1239.30\] For monthly compounded interest: \[A_{\text{monthly}} = 900 \cdot \left(1 + \frac{0.029}{12}\right)^{12 \cdot 13} = 900 \cdot (1.0024167)^{156} ≈ $1266.52\] Now, calculate the difference: \[\text{Difference} = A_{\text{monthly}} - A_{\text{simple}} = $1266.52 - $1239.30 ≈ $27.22\] So, the difference in the total account balances if simple interest is applied compared to monthly compounded interest is approximately $27.22.

SmoothCriminal:

Essentially the same thing as the previous problems.

Phantomdex:

I uh got the same question and did that answer the previous time and it said it was wrong-

SmoothCriminal:

@phantomdex wrote:
I uh got the same question and did that answer the previous time and it said it was wrong-
The one I just did it said was wrong? 😲

Phantomdex:

Yup

Phantomdex:

Or at least on the previous submission I did it was so-

Phantomdex:

Got another answer for me here?

SmoothCriminal:

@phantomdex wrote:
Got another answer for me here?
Yes, has to be $972.22 if not $27.22

Phantomdex:

I hope most of these are right or I will cry, because I have a borderline D in this class.

Phantomdex:

I got a 57

SmoothCriminal:

@phantomdex wrote:
I hope most of these are right or I will cry, because I have a borderline D in this class.
These are right, I next to never get these wrong. My teacher will smack me in my face if I get rusty already. I got a test next week.

SmoothCriminal:

@phantomdex wrote:
I got a 57
57 wasn't an answer choice tf-

Phantomdex:

No no the assignment was graded as a 57.

SmoothCriminal:

@phantomdex wrote:
No no the assignment was graded as a 57.
OHHHHH

SmoothCriminal:

@smoothcriminal wrote:
@phantomdex wrote:
No no the assignment was graded as a 57.
OHHHHH
Oh DAMN 😨😐

Phantomdex:

Yeah my mums gonna murder me again whew

SmoothCriminal:

@phantomdex wrote:
Yeah my mums gonna murder me again whew
Are you talking about THIS test was a 57?

SmoothCriminal:

@smoothcriminal wrote:
@phantomdex wrote:
Yeah my mums gonna murder me again whew
Are you talking about THIS test was a 57?
I mean assignment.

Phantomdex:

Yeah and it was my last submission sooo-

Phantomdex:

Not your fault my cousin had no clue what he was doing for two of em so

SmoothCriminal:

@phantomdex wrote:
Not your fault my cousin had no clue what he was doing for two of em so
Damn, you should have thought of me sooner 😭

Phantomdex:

@smoothcriminal wrote:
@phantomdex wrote:
Not your fault my cousin had no clue what he was doing for two of em so
Damn, you should have thought of me sooner 😭
I have more assignments so-

SmoothCriminal:

@phantomdex wrote:
@smoothcriminal wrote:
@phantomdex wrote:
Not your fault my cousin had no clue what he was doing for two of em so
Damn, you should have thought of me sooner 😭
I have more assignments so-
Well, I'll be here all night if you need me.

Phantomdex:

A small business owner has decided to invest some of the business profits to save for retirement. A total of $25,000 is invested with a continuously compounded annual interest rate of 2.9%. What is the total account balance after 20 years? $39,500 $46,450.96 $44,557.63 $44,650.96

SmoothCriminal:

@phantomdex wrote:
A small business owner has decided to invest some of the business profits to save for retirement. A total of $25,000 is invested with a continuously compounded annual interest rate of 2.9%. What is the total account balance after 20 years? $39,500 $46,450.96 $44,557.63 $44,650.96
D. $44,650.96.

Phantomdex:

A public relations writer received a promotion that came with a $5,000 salary increase. The writer wants to invest the money in a savings account with an annual interest rate of 5.4% for 10 years. How much interest would the account earn with simple interest and with monthly compounded interest? The account will earn $2,700 in simple interest and $3,569.65 in compound interest. The account will earn $3,569.65 in simple interest and $2,700 in compound interest. The account will earn $2,070 in simple interest and $3,580.03 in compound interest. The account will earn $3,580.03 in simple interest and $2,070 in compound interest.

SmoothCriminal:

@phantomdex wrote:
A public relations writer received a promotion that came with a $5,000 salary increase. The writer wants to invest the money in a savings account with an annual interest rate of 5.4% for 10 years. How much interest would the account earn with simple interest and with monthly compounded interest? The account will earn $2,700 in simple interest and $3,569.65 in compound interest. The account will earn $3,569.65 in simple interest and $2,700 in compound interest. The account will earn $2,070 in simple interest and $3,580.03 in compound interest. The account will earn $3,580.03 in simple interest and $2,070 in compound interest.
Are there answer choices or nah?

Phantomdex:

the four ones that are separated from the top text-

SmoothCriminal:

@phantomdex wrote:
the four ones that are separated from the top text-
Oh I missed that 😩

SmoothCriminal:

@phantomdex wrote:
the four ones that are separated from the top text-
A. The account will earn $2,700 in simple interest and $3,569.65 in compound interest.

Phantomdex:

A recent high school graduate is taking out a $2,500 loan to pay for a technical education program. The loan will accrue simple interest until repayment begins in 2 years. If the graduate wants to cap the accrued interest at $150, what interest rate is needed? 30% 5.7% 3% 0.3%

SmoothCriminal:

@phantomdex wrote:
A recent high school graduate is taking out a $2,500 loan to pay for a technical education program. The loan will accrue simple interest until repayment begins in 2 years. If the graduate wants to cap the accrued interest at $150, what interest rate is needed? 30% 5.7% 3% 0.3%
D. 0.3%

Phantomdex:

The equation A = (400)(0.053)(5) + (400) represents the amount of money earned on a simple interest savings account. What does the value 400 represent? Time Principal amount Interest rate Final amount

SmoothCriminal:

@phantomdex wrote:
The equation A = (400)(0.053)(5) + (400) represents the amount of money earned on a simple interest savings account. What does the value 400 represent? Time Principal amount Interest rate Final amount
I believe it is B. Principal amount

Phantomdex:

A wedding photographer has been saving money to use for travel during retirement. The photographer plans to invest in a continuously compounded savings account with an annual interest rate of 4.3%. If the photographer wants to have a total account balance of $40,000 after 20 years, what principal amount must initially be invested? $23,073.52 $26,020.36 $13,979.64 $16,926.48